A student uses a pH meter to measure the pH of a 2.0 M solution of ammonium hydroxide. The pH meter reads 12.94. What is the pOH of the solution?
1
Expert's answer
2020-12-01T09:13:53-0500

pH+pOH=14

pOH =14 - pH = 14 - 12.94 = 1.06